If the volume of a cone is 12pi cubic feet, write the formula for the height (h) of this cone in terms of its radius (r).
Answer by mananth(16949)


V=1/3 * pi * r^2 h
12 *pi =1/3 * pi* r^2 *h
12*3/r^2=h
36/r^2=h